The tail of a kite is 1.5ft plus twice the length of the kite. Together, the kite and tail are 15ft 6in long. What is the length of the tail?
Answer by RAY100(1637)

Let k = length of kite
.
1.5 +2k = lg tail
.
Total lg = k + 1.5 +2k = 15' 6" = 186 "
.
3k = 186 - 1.5 ft = 186 -18 = 168 "
.
k= 168 /3 = 56 "
.
check, overall = 56 + 18 + 112 = 186 = 15'6",,,,ok
.